Find two consecutive integers such that the sum of 2 times the first integer and 4 times second integer is 46.
Answer by checkley75(3666)

the two numbers are x & (x+1)
2x+4(x+1)=46
2x+4x+4=46
6x=46-4
6x=42
x=42/6
x=7 answer for the first number &
7+1=8 for the second number
proof
2*7+4*8=46
14+32=46
46=46